True or false:


The binomial coefficient "\displaystyle{{-1/2}\choose{5}}" is equal to "\dfrac{63}{256}"

​.


1
Expert's answer
2021-09-22T17:52:29-0400

By definition, the binomial coefficient "\binom nk" is given as follows:


"\binom nk = \frac{n^{\underline{k}}}{k!} = \frac{n(n-1)(n-2)\cdots(n-(k-1))}{k(k-1)(k-2)\cdots 1}"

Substituting "k = 5" and "n = -1\/2", get:

"\binom {-1/2}5 =\frac{(-1/2)^{\underline{5}}}{5!} =\\
= \frac{\left( -\dfrac12 \right)\left( -\dfrac12-1 \right)\left( -\dfrac12-2 \right)\left( -\dfrac12-3 \right)\left( -\dfrac12 -4\right)}{120} =\\
= -\dfrac{63}{256}"

Answer. False.